# Breaking the Orthogonality Barrier in Quantum LDPC Codes

### v2

@@ Abstract
Classical low-density parity-check (LDPC) codes are a widely deployed and well-established technology, forming the backbone of modern communication and storage systems. It is well known that, in this classical setting, increasing the girth of the Tanner graph while maintaining regular degree distributions leads simultaneously to good belief-propagation (BP) decoding performance and large minimum distance. In the quantum setting, however, this principle does not directly apply because quantum LDPC codes must satisfy additional orthogonality constraints between their parity-check matrices. When one enforces both orthogonality and regularity in a straightforward manner, the girth is typically reduced and the minimum distance becomes structurally upper bounded.
In this work, we overcome this limitation by using permutation matrices with controlled commutativity and by restricting the orthogonality constraints to only the necessary parts of the construction, while preserving regular check-matrix structures. This design breaks the conventional trade-off between orthogonality, regularity, girth, and minimum distance, allowing us to construct quantum LDPC codes with large girth and without the usual distance upper bounds. As a concrete demonstration, we construct a girth-8, (3,12)-regular $9216,4612,\leq 48]]$ quantum LDPC code and show that, under BP decoding combined with a low-complexity post-processing algorithm, it achieves a frame error rate as low as $10^{-8}$ on the depolarizing channel with error probability $4\%$.
@@ 1 Introduction
In this work, we first discuss the general mechanism by which row deletion can degrade the minimum distance of CSS codes. Next, for generalized Hagiwara–Imai codes [14], we establish a code-design method that controls commutativity of submatrices to avoid distance degradation due to row deletion. More concretely, using the algebraic structure of APM–LDPC codes, we guarantee orthogonality only on the active part, introduce a set of row pairs for which commutation is required, and localize the commutation condition. We further control the nonzero pattern of the interaction matrix so that deleted rows do not materialize as low-weight logical operators, and present a sequential construction algorithm based on the APM composition rule. As a concrete demonstration, we construct a (3,12)-regular $9216,4612,\leq 48]]$ quantum LDPC code and show that, under BP decoding combined with a low-complexity post-processing algorithm, it achieves a frame error rate (FER) as low as $10^{-8}$ on the depolarizing channel with error probability $4\%$. We also compare constructions with and without commutation control to demonstrate suppression of low-weight logical operators.
@@ 2 Problem setting: low-weight logical operators induced by check-row removal
We abstract the conventional CSS–LDPC construction used in this paper, based on [18, 10]. We first construct two orthogonal square (block-circulant) mother matrices $\hat{H}_{X},\hat{H}_{Z}$. The CSS code defined by these mother matrices typically has rate zero, so to adjust the rate we delete internal rows and use the remaining rows as the active matrices $H_{X}$ and $H_{Z}$. Namely,
@@
However, in this construction, mother orthogonality imposes strong constraints on the latent rows and can degrade the minimum distance. The minimum distance is defined as $d_{\min}=\min\{d_{X},d_{Z}\}$ with
@@
In this work we also obtain the active matrices $H_{X},H_{Z}$ by deleting rows from the mother matrices, but we design them so that low-weight rows in the latent matrices are not orthogonal to $H_{Z}$ or $H_{X}$, respectively. That is, for low-weight $\boldsymbol{x}\in\operatorname{Row}(\tilde{H}_{X})$ we enforce $H_{Z}\boldsymbol{x}^{\mathsf{T}}\neq 0$, and for low-weight $\boldsymbol{z}\in\operatorname{Row}(\tilde{H}_{Z})$ we enforce $H_{X}\boldsymbol{z}^{\mathsf{T}}\neq 0$, so that in particular $\operatorname{Row}(\tilde{H}_{X})\not\subset C_{Z}$ and $\operatorname{Row}(\tilde{H}_{Z})\not\subset C_{X}$. The same argument applies not only to individual latent rows but also to low-weight linear combinations.
Accordingly, we define latent-based upper bounds for the $X$ and $Z$ distances as
$d_{X}^{(\mathrm{lat})}$ $:=\min\{\operatorname{wt}(\boldsymbol{x}):\boldsymbol{x}\in\operatorname{Row}(\tilde{H}_{X})\cap C_{Z}\setminus C_{X}^{\perp}\},$ (5)
$=\min\{\operatorname{wt}(\boldsymbol{x}):\boldsymbol{x}\in C_{Z},\ \boldsymbol{x}=(\tilde{H}_{X})^{\mathsf{T}}\boldsymbol{u},\ \text{for some }\boldsymbol{u},\ \boldsymbol{x}\notin C_{X}^{\perp}\},$ (6)
$d_{Z}^{(\mathrm{lat})}$ $:=\min\{\operatorname{wt}(\boldsymbol{z}):\boldsymbol{z}\in\operatorname{Row}(\tilde{H}_{Z})\cap C_{X}\setminus C_{Z}^{\perp}\},$ (7)
$=\min\{\operatorname{wt}(\boldsymbol{z}):\boldsymbol{z}\in C_{X},\ \boldsymbol{z}=(\tilde{H}_{Z})^{\mathsf{T}}\boldsymbol{v},\ \text{for some }\boldsymbol{v},\ \boldsymbol{z}\notin C_{Z}^{\perp}\}$ (8)
The proposed design aims to make these bounds as large as possible. In [12], binary submatrices corresponding to codewords are lifted to a nonbinary field so that they no longer represent codewords; this can be viewed as increasing these latent-based bounds via nonbinary lifting.
@@ 3 Method of matrix design
$H_{X}H_{Z}^{\mathsf{T}}=0,\qquad H_{X}(\tilde{H}_{Z})^{\mathsf{T}}\neq 0,\qquad H_{Z}(\tilde{H}_{X})^{\mathsf{T}}\neq 0$ (9)
and $d_{X}^{(\mathrm{lat})},d_{Z}^{(\mathrm{lat})}$ are large. We define generalized Hagiwara–Imai codes, i.e., mother matrix pairs with block-circulant structure, and develop a general theory for imposing only the active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$ between the active matrices $H_{X}$ and $H_{Z}$.
A protograph LDPC code with column weight $J$ and row weight $L$ is defined by a parity-check matrix consisting of $J\times L$ permutation matrices of size $P$ [30]. For a permutation $f:[P]\to[P]$, the corresponding $P\times P$ permutation matrix $F=P(f)$ is defined by $F_{x,y}=1\;\Longleftrightarrow\;f(x)=y.$ Each row and column has exactly one 1, so the mother matrix is sparse while its block structure is compact.
We introduce generalized Hagiwara–Imai codes [14] as protograph generalizations of the Hagiwara–Imai quasi-cyclic CSS codes [10]. For permutation matrices $F_{i},G_{i}\ (i\in[L_{2}])$ of size $P$, define the mother matrices $\hat{H}_{X},\hat{H}_{Z}$ of size $(L_{2}P)\times(LP)$ by
$(\hat{H}_{X})_{i,j}=F_{j-i},\quad(\hat{H}_{X})_{i,L_{2}+j}=G_{j-i},$ (10)
$(\hat{H}_{Z})_{i,j}=(G_{i-j})^{\mathsf{T}},\quad(\hat{H}_{Z})_{i,L_{2}+j}=(F_{i-j})^{\mathsf{T}}.$ (11)
All subscripts are taken modulo $L_{2}$. Throughout, indices in $[L/2]$ are taken modulo $L/2$, i.e., we identify $[L/2]$ with $\mathbb{Z}_{L/2}$. A sufficient condition for orthogonality is given below. The active matrices $H_{X}$ and $H_{Z}$ consist of the top $J$ block rows. Conventional designs impose commutativity of $F_{i}$ and $G_{j}$ so that both $H_{X},H_{Z}$ and $\hat{H}_{X},\hat{H}_{Z}$ are orthogonal. The top $J$ block rows of the mother matrices are used as the active matrices $H_{X}$ and $H_{Z}$. Our interest is to satisfy active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$ while not imposing mother-matrix orthogonality.
Each block row of $\hat{H}_{X}$ is a cyclic shift of $(F_{0},F_{1},\dots,F_{L/2-1})$ in the left half and $(G_{0},G_{1},\dots,G_{L/2-1})$ in the right half. Hence blocks depend only on the difference (ji)(j-i), and the product of mother matrices depends only on the difference. For any $i,k\in[L_{2}]$, the $(i,k)$ block is
$\bigl(\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}\bigr)_{i,k}=\sum_{u=0}^{L_{2}-1}\Bigl(F_{u}G_{k-u}+G_{k-u}F_{u}\Bigr)=:\Psi_{r}$ (12)
which depends only on $r=(k-i)\bmod\,L_{2}$. For $r\in[L/2]$, if $F_{u}$ and $G_{r-u}$ commute for all $u\in[L/2]$, then $\Psi_{r}=0$.
@@ 3.3 Upper bound on girth
Finally, we recall a structural limitation that applies when commutativity holds on all required differences.
@@
Let $L$ be even and assume $2\leq J\leq L/2$. If $F_{u}$ and $G_{r-u}$ commute for all $r\in\Delta$ and all $u\in[L/2]$, then the Tanner graphs of $H_{X}$ and $H_{Z}$ each contain an 8-cycle.
@@
Since $J\geq 2$, we have $0,1\in\Delta$. Also $L/2\geq 2$, so there exist $i\neq i^{\prime}$ in $[L/2]$. Let $r=0$, $s=1$, and set j=rij=r-i, $j^{\prime}=s-i$ (indices modulo $L/2$). By definition, $(i,j),(i,j^{\prime})\in\Gamma_{r}\cup\Gamma_{s}$, and similarly $(i^{\prime},j),(i^{\prime},j^{\prime})\in\Gamma_{r}\cup\Gamma_{s}$. By assumption the corresponding $F$ and $G$ commute, so the 8-cycle word composed of the two $F$ columns and two $G$ columns is
$W=F_{i}G_{j}^{-1}G_{j^{\prime}}F_{i}^{-1}F_{i^{\prime}}G_{j^{\prime}}^{-1}G_{j}F_{i^{\prime}}^{-1}$ (14)
Using commutativity to reorder gives $W=(F_{i}F_{i}^{-1})(F_{i^{\prime}}F_{i^{\prime}}^{-1})(G_{j}^{-1}G_{j^{\prime}})(G_{j^{\prime}}^{-1}G_{j})=I$, so an 8-cycle necessarily exists in the Tanner graphs of $H_{X}$ and $H_{Z}$. ∎
@@ 4 Method of Matrix Construction
    At least one $(i,j)\in[L/2]^{2}\setminus\Delta$ is non-commuting.
@@ 5 Method of Decoding
Our experiments use BP decoding [18, 23] as the baseline. We first recall a basic principle: if we knew the support where the BP estimate differs from the true error, decoding would reduce to a linear system on the corresponding columns, and a unique solution would imply successful recovery. For example, if the $X$-side mismatch support is $E$, then $H_{Z}(:,E)\,\boldsymbol{e}_{E}=\boldsymbol{r}_{X}$; if $H_{Z}(:,E)$ has full column rank, the solution is unique and yields the correct correction. The same applies to the $Z$ side with $H_{X}$ and $\boldsymbol{r}_{Z}$.
@@
Empirically, BP stalls or slowly oscillates roughly once per $10^{5}$ trials, and we trigger post-processing in such cases. We invoke three post-processing methods: ETS-based post-processing, a correction based on flip history, and an OSD-based method. Below we briefly describe the procedure and the criteria. In post-processing, the residual syndrome $\boldsymbol{r}$ is $\boldsymbol{r}_{X}$ (with $H=H_{Z}$) or $\boldsymbol{r}_{Z}$ (with $H=H_{X}$), depending on the side.
@@ 5.1 Flip history decoding
Flip history decoding (FHD) forms a candidate set $F$ from variables that flipped during BP and seeks a correction that cancels the residual syndrome on the local submatrix. This local correction is a heuristic to resolve stalls caused by trapping structures [27]. Related local-flip decoders include weighted bit-flipping [34]. Concretely, let $N(F)$ be the set of check rows adjacent to $F$, and solve the GF(2) linear system
$H_{N(F),F}\,\boldsymbol{\delta}=\boldsymbol{r}_{N(F)}.$
If the system is solvable and the solution is unique, we flip the bits indicated by $\boldsymbol{\delta}$.
5.2 OSD
OSD [4] uses the least reliable variables from BP to form a column set $K$ and solves
$H_{:,K}\,\boldsymbol{x}=\boldsymbol{r}$
using all rows. The method is applied independently to the $X$ and $Z$ sides. In the implementation, we perform a binary search on $|K|$ to find the minimum $K$ that is solvable, then take the largest $K$ within the range that preserves uniqueness. The correction is applied only when the solution weight is below a threshold.
5.3 Elementary Trapping Set (ETS)-based post-processing
ETS-based post-processing targets elementary trapping sets with $b=2$ unsatisfied checks, which are among the most harmful structures that cause BP to stall. We precompute a library of ETSs from the Tanner graph (e.g., $(6,2)$, $(8,2)$, $(10,2)$) and store each entry as a variable set $V$ and its odd check pair $(c_{0},c_{1})$. During decoding, we run ETS-based post-processing only when the number of unsatisfied checks is small (we skip post-processing when the number is not small) and the residual syndrome has exactly two unsatisfied checks.
Given the current estimate, let $\boldsymbol{r}$ be the residual syndrome and $U=\{c_{0},c_{1}\}$ be its unsatisfied check pair. We scan ETS entries whose odd checks match $U$. For each candidate ETS with variable set $V$, we build the local check set $N(V)$ and reject it if $\boldsymbol{r}$ has nonzero entries in $N(V)\setminus U$. Otherwise we solve the local system
$H_{N(V),V}\,\boldsymbol{\delta}=\boldsymbol{r}_{N(V)}$
over $\mathbb{F}_{2}$ using Gaussian elimination. If solvable, we flip variables in $V$ according to $\boldsymbol{\delta}$ and accept the correction only when the syndrome matches the target. We try ETS sizes in increasing order and apply the first successful correction, independently for the $X$ and $Z$ sides.

### v3

@@ Abstract
Classical low-density parity-check (LDPC) codes are a widely deployed and well-established technology, forming the backbone of modern communication and storage systems. It is well known that, in this classical setting, increasing the girth of the Tanner graph while maintaining regular degree distributions leads simultaneously to good belief-propagation (BP) decoding performance and large minimum distance. In the quantum setting, however, this principle does not directly apply because quantum LDPC codes must satisfy additional orthogonality constraints between their parity-check matrices. When one enforces both orthogonality and regularity in a straightforward manner, the girth is typically reduced and the minimum distance becomes structurally upper bounded. In this work, we overcome this limitation by using permutation matrices with controlled commutativity and by restricting the orthogonality constraints to only the active part of the construction, while preserving regular check-matrix structures. This design circumvents conventional structural distance limitations induced by parent-matrix orthogonality, and enables the construction of quantum LDPC codes with large girth while avoiding latent low-weight logical operators. As a concrete demonstration, we construct a girth-8, (3,12)-regular $9216,4612,\leq 48]]$ quantum LDPC code and show that, under BP decoding combined with a low-complexity post-processing algorithm, it achieves a frame error rate as low as $10^{-8}$ on the depolarizing channel with error probability $4\%$.
@@ 1 Introduction
In this work, we first discuss the general mechanism by which row deletion can degrade the minimum distance of CSS codes. Next, for generalized Hagiwara–Imai codes [14], we establish a code-design method that controls commutativity of submatrices to avoid distance degradation due to row deletion. More concretely, using the algebraic structure of APM–LDPC codes, we guarantee orthogonality only on the active part, introduce a set of row pairs for which commutation is required, and localize the commutation condition. We further control the nonzero pattern of the interaction matrix so that deleted rows do not materialize as low-weight logical operators, and present a sequential construction algorithm based on the APM composition rule. We also give a girth upper bound under the required commutativity and explicitly construct a code that attains this bound.
As a concrete demonstration, we construct a (3,12)-regular $9216,4612,\leq 48]]$ quantum LDPC code and show that, under BP decoding combined with a low-complexity post-processing algorithm, it achieves a frame error rate (FER) as low as $10^{-8}$ on the depolarizing channel with error probability $4\%$. We also compare constructions with and without commutation control to demonstrate suppression of low-weight logical operators.
@@ 2 Problem setting: low-weight logical operators induced by check-row removal
We abstract the conventional CSS–LDPC construction used in this paper, based on [18, 10]. We first construct two orthogonal square (block-circulant) parent matrices $\hat{H}_{X},\hat{H}_{Z}$. The CSS code defined by these parent matrices typically has rate zero, so to adjust the rate we delete internal rows and use the remaining rows as the active matrices $H_{X}$ and $H_{Z}$. Namely,
@@
However, in this construction, parent orthogonality imposes strong constraints on the latent rows and can degrade the minimum distance. The minimum distance is defined as $d_{\min}=\min\{d_{X},d_{Z}\}$ with
@@
In this work we also obtain the active matrices $H_{X},H_{Z}$ by deleting rows from the parent matrices, but we design them so that low-weight rows in the latent matrices are not orthogonal to $H_{Z}$ or $H_{X}$, respectively. That is, for low-weight $\boldsymbol{x}\in\operatorname{Row}(\tilde{H}_{X})$ we enforce $H_{Z}\boldsymbol{x}^{\mathsf{T}}\neq 0$, and for low-weight $\boldsymbol{z}\in\operatorname{Row}(\tilde{H}_{Z})$ we enforce $H_{X}\boldsymbol{z}^{\mathsf{T}}\neq 0$, so that in particular $\operatorname{Row}(\tilde{H}_{X})\not\subset C_{Z}$ and $\operatorname{Row}(\tilde{H}_{Z})\not\subset C_{X}$. The same argument applies not only to individual latent rows but also to low-weight linear combinations.
Accordingly, we define latent-based distance for the $X$ and $Z$ distances as
$d_{X}^{(\mathrm{lat})}$ $:=\min\{\operatorname{wt}(\boldsymbol{x}):\boldsymbol{x}\in(C_{Z}\cap\operatorname{Row}(\tilde{H}_{X}))\setminus C_{X}^{\perp}\},$ (5)
$d_{Z}^{(\mathrm{lat})}$ $:=\min\{\operatorname{wt}(\boldsymbol{z}):\boldsymbol{z}\in(C_{X}\cap\operatorname{Row}(\tilde{H}_{Z}))\setminus C_{Z}^{\perp}\}.$ (6)
These quantities upper bound $d_{\min}$. The proposed design aims to make these bounds as large as possible. In [12], binary submatrices corresponding to codewords are lifted to a nonbinary field so that they no longer represent codewords; this can be viewed as increasing these latent-based distance via nonbinary lifting.
@@ 3 Method of matrix design
$H_{X}H_{Z}^{\mathsf{T}}=0,\qquad H_{X}(\tilde{H}_{Z})^{\mathsf{T}}\neq 0,\qquad H_{Z}(\tilde{H}_{X})^{\mathsf{T}}\neq 0,$ (7)
and, more precisely, such that $d_{X}^{(\mathrm{lat})}$ and $d_{Z}^{(\mathrm{lat})}$ are large. We define generalized Hagiwara–Imai codes, i.e., parent matrix pairs with block-circulant structure, and develop a general theory for imposing only the active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$ between the active matrices $H_{X}$ and $H_{Z}$.
A protograph LDPC code with column weight $J$ and row weight $L$ is defined by a parity-check matrix consisting of $J\times L$ permutation matrices of size $P$ [31]. For a permutation $f:[P]\to[P]$, the corresponding $P\times P$ permutation matrix $F=P(f)$ is defined by $F_{x,y}=1\text{ if and only if }f(x)=y.$ Each row and column has exactly one 1, so the parent matrix is sparse while its block structure is compact.
We introduce generalized Hagiwara–Imai codes [14] as protograph generalizations of the Hagiwara–Imai quasi-cyclic CSS codes [10]. For permutation matrices $F_{i},G_{i}\ (i\in[L/2])$ of size $P$, define the parent matrices $\hat{H}_{X},\hat{H}_{Z}$ of size $(LP/2)\times(LP)$ by
$(\hat{H}_{X})_{i,j}=F_{j-i},\quad(\hat{H}_{X})_{i,L/2+j}=G_{j-i},$ (8)
$(\hat{H}_{Z})_{i,j}=(G_{i-j})^{\mathsf{T}},\quad(\hat{H}_{Z})_{i,L/2+j}=(F_{i-j})^{\mathsf{T}}.$ (9)
Throughout, indices in $[L/2]$ are taken modulo $L/2$, i.e., we identify $[L/2]$ with $\mathbb{Z}_{L/2}$. A sufficient condition for orthogonality is given below. The active matrices $H_{X}$ and $H_{Z}$ consist of the top $J$ block rows. Conventional designs [14, 12] impose commutativity of $F_{i}$ and $G_{j}$ so that both $H_{X},H_{Z}$ and $\hat{H}_{X},\hat{H}_{Z}$ are orthogonal. The top $J$ block rows of the parent matrices are used as the active matrices $H_{X}$ and $H_{Z}$. Our interest is to satisfy active orthogonality $H_{X}H_{Z}^{\mathsf{T}}=0$ while not imposing parent-matrix orthogonality.
This focus on active orthogonality is motivated by the desire to keep the latent part non-orthogonal to the active part, so that low-weight latent combinations do not automatically become logical operators. The block-circulant structure makes $\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}$ depend only on differences, and the resulting interaction matrices $\Psi_{r}$ allow us to localize commutativity constraints to a small difference set.
Each block row of $\hat{H}_{X}$ is a cyclic shift of $(F_{0},F_{1},\dots,F_{L/2-1})$ in the left half and $(G_{0},G_{1},\dots,G_{L/2-1})$ in the right half. Hence blocks depend only on the difference (ji)(j-i), and the product of parent matrices depends only on the difference. For any $i,k\in[L/2]$, the $(i,k)$ block is
$\bigl(\hat{H}_{X}(\hat{H}_{Z})^{\mathsf{T}}\bigr)_{i,k}=\sum_{u=0}^{L/2-1}\Bigl(F_{u}G_{r-u}+G_{r-u}F_{u}\Bigr)=:\Psi_{r}$ (10)
which depends only on $r:=(k-i)\bmod\,L/2$. For $r\in[L/2]$, if $F_{u}$ and $G_{r-u}$ commute for all $u\in[L/2]$, then $\Psi_{r}=0$.
@@ 3.2 Necessary condition for latent non-orthogonality
Intuitively, when $L<4J$ the difference set $\Delta$ covers all residues, so enforcing active orthogonality forces $\Psi_{r}=0$ for every $r$. This annihilates all mixed active–latent blocks and makes latent rows orthogonal, which is exactly what we aim to avoid.
@@ 3.3 Upper bound on girth
Finally, we recall a structural limitation that applies to the active matrices when commutativity holds on all required differences.
@@
Let $L$ be even and assume $3\leq J\leq L/2$ with the standard active choice (top $J$ block rows). If $F_{u}$ and $G_{r-u}$ commute for all $r\in\Delta$ and all $u\in[L/2]$, then the Tanner graphs of the active matrices $H_{X}$ and $H_{Z}$ each contain an 8-cycle.
@@
Since $J\geq 3$, the active block rows include $0,1,2$. Choose $i,j\in[L/2]$ so that $i+j,i+j-1,i+j-2\in\Delta$ (e.g., $i=j=0$). Consider the block positions on the active rows
[(0,i),(0,L/2+j),(1,L/2+j),(1,i+1),(2,i+1),(2,L/2+j+1),(1,L/2+j+1),(1,i)],[(0,i),(0,L/2+j),(1,L/2+j),(1,i+1),(2,i+1),(2,L/2+j+1),(1,L/2+j+1),(1,i)], (12)
where indices are modulo $L/2$. The corresponding cycle product, i.e., the product of permutation matrices encountered along this closed walk, is
$W=F_{i}G_{j}^{-1}G_{j-1}F_{i}^{-1}F_{i-1}G_{j-1}^{-1}G_{j}F_{i-1}^{-1}$ (13)
and the pairs $(i,j)$, (i,j1)(i,j-1), (i1,j)(i-1,j), and (i1,j1)(i-1,j-1) correspond to the differences $i+j$, i+j1i+j-1, and i+j2i+j-2 in $\Delta$. By assumption the relevant $F$ and $G$ commute, so $W=I$ and an 8-cycle exists in the active Tanner graphs. ∎
@@ 4 Method of Matrix Construction
    At least one $(i,j)\in[L/2]^{2}\setminus\Gamma$ is non-commuting.
@@ 5 Method of Decoding
Our experiments use BP decoding [18, 24] as the baseline. We first recall a basic principle: if we knew the support where the BP estimate differs from the true error, decoding would reduce to a linear system on the corresponding columns, and a unique solution would imply successful recovery. We write $H[A,B]$ for the submatrix with row set $A$ and column set $B$, and for a vector $\boldsymbol{v}$ we write $\boldsymbol{v}[A]$ for its restriction to $A$. For example, if the $X$-side mismatch support is $E$, then with $A$ the full check set, $H_{Z}[A,E]\,\boldsymbol{e}[E]=\boldsymbol{r}_{X}[A]$; if $H_{Z}[A,E]$ has full column rank, the solution is unique and yields the correct correction. The same applies to the $Z$ side with $H_{X}$ and $\boldsymbol{r}_{Z}$.
@@
Empirically, BP stalls or slowly oscillates roughly once per $10^{5}$ trials (frames), and we trigger post-processing in such cases, focusing on instances with a small number of unsatisfied checks; in our experiments we activate it only when the number of unsatisfied checks is at most 20.
We invoke three post-processing methods: ETS-based post-processing, a correction based on flip history, and an OSD-based method. Below we briefly describe the procedure and the criteria. In post-processing, the residual syndrome $\boldsymbol{r}$ is $\boldsymbol{r}_{X}$ (with $H=H_{Z}$) or $\boldsymbol{r}_{Z}$ (with $H=H_{X}$), depending on the side.
[…]
@@ 5.1 Flip history decoding
Flip history decoding (FHD) uses the flip history to form a candidate support $\hat{E}=F$, where $F$ is the set of variables that flipped during BP, and then applies the local linear solve on $N(F)$. This local correction is a heuristic to resolve stalls caused by trapping structures [28]. Related local-flip decoders include weighted bit-flipping [35].
5.2 Ordered statistics decoding
Ordered statistics decoding (OSD) [4] uses the least reliable variables from BP to form a candidate support $\hat{E}=K$. In the implementation, we perform a binary search on $|K|$ to find the minimum $K$ that is solvable, then take the largest $K$ within the range that preserves uniqueness.
5.3 Trapping set based post-processing
An $(a,b)$ elementary trapping set is a variable-node set $V$ with $|V|=a$ whose induced subgraph has exactly $b$ unsatisfied (odd-degree) check nodes; it is elementary if every check node in the induced subgraph has degree 1 or 2. The post-processing targets elementary trapping sets with $b=2$ unsatisfied checks, which are among the most harmful structures that cause BP to stall. We precompute a library of harmful trapping sets from the Tanner graph and store each entry as a variable set $V$ (used as $\hat{E}$) and its odd check pair $(c_{0},c_{1})$. During decoding, we run the post-processing only when the residual syndrome has exactly two unsatisfied checks.